# An Efficient Finite Difference Approximation via a Double Sample-Recycling Approach

### v1

@@ Abstract
Estimating stochastic gradients is pivotal in fields like service systems within operations research. The classical method for this estimation is the finite difference approximation, which entails generating samples at perturbed inputs. Nonetheless, practical challenges persist in determining the perturbation and obtaining an optimal finite difference estimator in the sense of possessing the smallest mean squared error (MSE). To tackle this problem, we propose a double sample-recycling approach in this paper. Firstly, pilot samples are recycled to estimate the optimal perturbation. Secondly, recycling these pilot samples again and generating new samples at the estimated perturbation, lead to an efficient finite difference estimator. We analyze its bias, variance and MSE. Our analyses demonstrate a reduction in asymptotic variance, and in some cases, a decrease in asymptotic bias, compared to the optimal finite difference estimator. Therefore, our proposed estimator consistently coincides with, or even outperforms the optimal finite difference estimator. In numerical experiments, we apply the estimator in several examples, and numerical results demonstrate its robustness, as well as coincidence with the theory presented, especially in the case of small sample sizes.
Key words: Finite difference; Sample-recycling; Mean square error; Perturbation; Bootstrap.
@@ 1 Introduction
Gradient estimation involves estimating a function’s derivative. In realistic scenarios, the exact function expression is often unknown, and one can only approximate it through stochastic simulation. Stochastic gradient estimation finds significant applications in stochastic optimization, machine learning, finance, and operations research.
Stochastic gradient estimation has substantial applications in queuing theory, one of the most important fields of study in operations research (see Fu, 2006; Lam et al., 2022; Li and Lam, 2020). In this field, the primary objective function pertains to the performance of a queuing system, such as average waiting time or queue length. A stochastic service system can be designed and controlled more effectively by analyzing the probabilities involved in queueing and waiting phenomena. Queuing theory finds extensive applications in resource allocation management for manufacturing, product production, transportation, and inventory. Stochastic gradient estimation methods can be employed to estimate the gradient of the queuing model performance function to the model parameters. In this manner, efficient optimization of queueing systems can be achieved, thereby improving overall performance and enhancing the efficiency of operations.
The finite difference technique (see Asmussen and Glynn, 2007; L’Ecuyer, ; Zazanis and Suri, 1993) is a classic approach for estimating the gradient of a function. Forward, backward, and central finite difference (CFD) methods are examples of the finite difference methods. Finite difference estimation necessitates the creation of samples with varying input parameters. The perturbation size is determined by reducing the estimator’s mean squared error (MSE). In contrast, increasing the perturbation leads to both an increase in the bias and a decrease in the variance. To reduce the MSE, the bias and variance must be balanced. Without using common random numbers (CRNs), the optimal perturbation size of the CFD technique is $cn^{-1/6}$, corresponding to the optimal MSE of order $n^{-2/3}$, where $n$ is the number of different pairs in the simulation (Fox and Glynn, 1989). The ideal perturbation size in the forward or backward finite difference method is $c_{1}n^{-1/4}$, and the corresponding optimal MSE order is $n^{-1/2}$. Despite the fact that the choice of perturbation size is widely regarded as the heart of the finite difference technique in relevant studies, the perturbation selection approach has not been fully established. Although the order of the optimal perturbation corresponding to the various finite difference methods is known, the frontal constants, i.e., $c$ and $c_{1}$, remain unknown. Experiments demonstrate that changing the constant value has a significant influence on the MSE, especially for finite samples. However, estimating unknown constants is sometimes seen as more difficult since it requires more model information, such as higher-order derivatives and noise variance. These details are unknown, and even approximating them is a more challenging task than the gradient estimation itself. This work explores the CFD technique, with the goal of extracting this information and providing a more accurate CFD estimator.
The closest work to ours is Li and Lam (2020). They suggest a two-step technique that uses a few samples in the first stage to roughly estimate the third-order derivative of the function using regression and estimate the noise using the sample variance. After this, they obtained an estimator of the optimal perturbation. Then, they construct a finite difference estimator according to the estimated perturbation in the second stage. Li and Lam (2020) referred to their method as the estimation-minimization CFD (EM-CFD) algorithm. In order to more accurately estimate the unknown constants and effectively utilize the samples, we propose a double sample-recycling (DSR) approach to the finite difference estimator for stochastic gradient estimation in this paper. Our work also employs the two-stage strategy but from a sample-recycling angle. That is, we apply two sample-recycling techniques in the two stages to obtain accurate estimators of perturbation and finite difference estimator. Specifically, we make the following contributions.
  1. 1.
    We propose a sample-recycling method based on the bootstrap technique to estimate the constants. The first stage devotes to estimating the constants, primarily relying on the functional relationships between the bias and variance of the finite difference estimator and the perturbation. We apply the sample-recycling technique, specifically the bootstrap method, on a set of pilot samples to estimate the bias and variance. Then the constants can be estimated through regression, utilizing the established functional relationships. The sample-recycling technique not only saves samples but also effectively estimates the bias and variance, so leading to relatively accurate estimators of the constants and then the optimal perturbation.
  2. 2.
    We provide theoretical analyses of the asymptotic properties in the first stage. We show that the bias and variance estimated using the bootstrap method have the same asymptotic properties as the theoretical ones, thereby preserving the aforementioned functional relationships. This ensures the validity of our proposed estimators for the constants and the optimal perturbation. Furthermore, we provide the convergence rate of these estimators, serving as a guide for the implementation of the first stage.
  3. 3.
    We construct a DSR finite difference estimator and prove its theoretical properties In the second stage, we generate remaining samples at the estimated optimal perturbation. Then we combine these samples with the pilot ones to construct the finite difference estimator. Specifically, we recycle the pilot samples by transforming them to ones with the same bias and variance as the remaining samples. Our theoretical analysis reveals a reduction in the variance of the DSR estimator compared to the optimal finite difference estimator and, in some cases, a decrease in bias.
  4. 4.
    We propose an efficient algorithm for practical implementation. Theoretical properties provide a guiding framework for practical implementation. Based on this, we propose an efficient algorithm to derive a robust finite difference estimator in practice. In particular, we apply the weighted least squares to mitigate the heteroscedasticity in the regression.
  5. 5.
    We demonstrate the practical benefits of DSR method via numerical experiments. We apply the DSR method in several examples. The results obtained from these numerical examples demonstrate the good performance of the proposed DSR method in terms of estimators of the constants, perturbation, and the finite difference estimator. The numerical results indicate the advantage of the DSR method over other approaches, especially in the case of small sample sizes.
@@
The rest of this paper is organized as follows. In Section 2, we give some background on how to use the CFD method for gradient estimation. In Section 3, we introduce the bootstrap method, apply it to gradient estimation and provide theoretical guarantees for it. Section 4 guides the process of reusing samples and provides theoretical evidence for the effectiveness of this operation. For implementation, an efficient algorithm is proposed in Section 5 and in Section 6, we give some experiments to verify the theoretical results, followed by conclusions in Section 7. Proofs are provided in the appendix.
@@ 2 Background
In this section, we introduce the background of the finite difference method for stochastic gradient estimation. More details can be found in Asmussen and Glynn (2007), Fu (2006) and Glasserman (2013). Consider the models that depend on a single parameter $\theta$, where $\theta$ varies within some range $\Theta\subset\mathbb{R}$. Denote that $\alpha(\cdot)$ is a performance measure of interest and assume that we can only estimate it by simulation. Within the simulation for any chosen $\theta\in\Theta$, each trail gives an unbiased but noisy estimate of $\alpha(\theta)$, denoted by $Y(\theta)$, i.e., $\alpha(\theta)=\mathbb{E}[Y(\theta)]$. Suppose that we do not use the CRNs, i.e., for $\theta_{1}\neq\theta_{2}\in\Theta$, $Y(\theta_{1})$ and $Y(\theta_{2})$ are independent. We would like to estimate the first-order derivative $\alpha^{\prime}(\theta_{0})$, where $\theta_{0}\in\Theta$ is the point of interest. To estimate $\alpha^{\prime}(\theta_{0})$, we prefer the CFD method over the forward (or backward) finite difference method because compared to the forward (and backward) finite difference method, the CFD method yields an improvement in the convergence rate of the bias. It is worth mentioning that the methods in Sections 3 and 4 are seamlessly adapted to the forward (and backward) finite difference scheme.
@@
Specifically, the CFD scheme specifies a perturbation parameter $h>0$. At the point $\theta_{0}+h$, the simulation is repeated independently, and $n$ independently and identically distributed (i.i.d.) observations $\{Y_{1}(\theta_{0}+h),...,Y_{n}(\theta_{0}+h)\}$ are obtained. Likewise, at $\theta_{0}-h$, $n$ i.i.d. observations $\{Y_{1}(\theta_{0}-h),...,Y_{n}(\theta_{0}-h)\}$ are generated. Using the sample average, we construct a CFD estimator of $\alpha^{\prime}(\theta_{0})$, denoted by
$\widehat{\Delta}_{n,h}\triangleq\frac{1}{n}\sum_{i=1}^{n}\Delta_{i}(h)=\frac{1}{n}\sum_{i=1}^{n}\frac{Y_{i}(\theta_{0}+h)-Y_{i}(\theta_{0}-h)}{2h}.$ (3)
For any $i=1,...,n$, $Y_{i}(\theta_{0}+h)$ and $Y_{i}(\theta_{0}-h)$ are the unbiased estimators of $\alpha(\theta_{0}+h)$ and $\alpha(\theta_{0}-h)$, respectively, so $\Delta_{i}(h)$ and $\widehat{\Delta}_{n,h}$ are both the unbiased estimators of $\widetilde{\alpha}_{h}^{\prime}(\theta_{0})$. Let $Z_{i}(h)$ denote the zero-mean error associated with $\Delta_{i}(h)$, i.e., $\Delta_{i}(h)=\widetilde{\alpha}_{h}^{\prime}(\theta_{0})+Z_{i}(h)$. Then, denote
$\bar{Z}_{n,h}=\frac{1}{n}\sum_{i=1}^{n}Z_{i}(h)$
be the zero-mean error associated with $\widehat{\Delta}_{n,h}$, i.e.,
@@
The selection of $h$ holds significant importance as it directly influences the quality of $\widehat{\Delta}_{n,h}$. If $h$ is too large, the expected value of $\widehat{\Delta}_{n,h}$, i.e., $\widetilde{\alpha}_{h}^{\prime}(\theta_{0})$ could deviate significantly from $\alpha^{\prime}(\theta_{0})$, resulting in an unacceptable bias. Conversely, if $h$ is too small, the variance of $\widehat{\Delta}_{n,h}$ will explode due to its position in the denominator. To choose a proper $h$, the challenge lies in striking a balance between mitigating the bias and restraining the variance of $\widehat{\Delta}_{n,h}$. To calculate the bias and variance elaborately, we make the following assumptions.
@@
The standard deviation of $Y(\theta)$, denoted by $\sigma(\theta)$, is continuous at $\theta_{0}$ and $\sigma(\theta)$ is bounded below away from 0.
@@
Subtraction eliminates $\alpha^{(2)}(\theta_{0})$, leaving
$\widetilde{\alpha}_{h}^{\prime}(\theta_{0})=\alpha^{\prime}(\theta_{0})+Bh^{2}+o(h^{2}),$
where $B=\alpha^{(3)}(\theta_{0})/6$. Therefore, the bias of $\widehat{\Delta}_{n,h}$ is
@@
It follows from (7) that $h$ controls a trade-off between the bias and variance, and thus we should choose an appropriate $h$ to minimize the MSE. According to the relationship between the arithmetic average and the geometric average, we get
$B^{2}h^{4}+\frac{\sigma^{2}(\theta_{0})}{2nh^{2}}\geq 3\left(\frac{B^{2}\sigma^{4}(\theta_{0})}{16n^{2}}\right)^{1/3},$
where the equality holds if and only if
@@
The optimal bias, variance and MSE are thus
@@
In order to obtain an estimator $\widehat{\Delta}_{n,h}$ with optimal MSE, it is important to determine the value of $h^{*}$ according to (8). Although it is established that $h^{*}=Cn^{-1/6}$ from (8), where $C=\left[\sigma^{2}(\theta_{0})/(4B^{2})\right]^{1/6}$, the constant $C$ is typically unknown, and involves two constants $\alpha^{(3)}(\theta_{0})$ and $\sigma^{2}(\theta_{0})$ that depend on model information. However, estimating these two constants is more challenging than estimating the gradient itself because they involve the higher-order derivatives of $\alpha(\theta)$. Furthermore, experimental findings indicate that assigning a perturbation with a significant deviation from the true one leads to substantial discrepancies in the MSE of the CFD estimator. These challenges motivate us to explore an efficient approach to estimate the constant $C$ so that the CFD estimator achieves its optimal MSE.
3 Sample Recycling for Perturbation Selection
In this section, we propose a sample recycling procedure aimed at estimating $C$, specifically $B$ and $\sigma^{2}(\theta_{0})$. Our inspiration stems from equations (5) and (6), which reveal that $\mathbb{E}\widehat{\Delta}_{n,h}$ and $\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]$ are linear w.r.t. $h^{2}$ and $1/h^{2}$, respectively. If we can accurately calculate $\mathbb{E}\widehat{\Delta}_{n,h}$ and $\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]$ for any given $h$, then choosing $h_{1}\neq h_{2}$ and solving the system of linear equations

### v4

# A Correlation-induced Finite Difference Estimator

@@ Abstract
Finite difference (FD) approximation is a classic approach to stochastic gradient estimation when only noisy function realizations are available. In this paper, we first provide a sample-driven method via the bootstrap technique to estimate the optimal perturbation, and then propose an efficient FD estimator based on correlated samples at the estimated optimal perturbation. Furthermore, theoretical analyses of both the perturbation estimator and the FD estimator reveal that, surprisingly, the correlation enables the proposed FD estimator to achieve a reduction in variance and, in some cases, a decrease in bias compared to the traditional optimal FD estimator. Numerical results confirm the efficiency of our estimators and align well with the theory presented, especially in scenarios with small sample sizes. Finally, we apply the estimator to solve derivative-free optimization (DFO) problems, and numerical studies show that DFO problems with 100 dimensions can be effectively solved.
Key words: finite difference; correlation; perturbation; mean squared error; DFO.
@@ 1 Introduction
Stochastic gradient estimation is an active field in simulation optimization and machine learning (Mohamed et al., 2020). In realistic scenarios, as the systems become more complex, their closed-form expressions are often unknown, and systems’ outputs can only be assessed using their stochastic realizations. Therefore, the derivative information about these systems is typically unavailable, characterizing this scenario as a black-box or zeroth-order model. In such settings, the FD approximation is a classic approach to gradient estimation (Asmussen and Glynn, 2007; L’Ecuyer, 1991; Zazanis and Suri, 1993).
FD approximation finds important applications in operations research and machine learning. A well-known application is sensitivity analysis for stochastic systems, such as queuing systems (Fu, 2015). In queueing systems, it is of interest that the sensitivity analysis for the average system with respect to (w.r.t.) the input parameters, such as ones in the arrival and service time distributions (Li and Lam, 2020; Lam et al., 2023).
Besides applications in sensitivity analysis, other important applications include solving the DFO problem where derivative information is unavailable. In practice, increasingly complex and diverse problems lack gradient information and only offer stochastic realizations of a function. Optimizing these problems has driven the development of DFO (Larson et al., 2019). For example, Google (Golovin et al., 2017) has established a Google-internal service called Google Vizier based on DFO methods for optimizing many machine learning methods and complex systems. In resent years, the use of FD estimators as surrogates for gradients in descent algorithms yields valid DFO algorithms, which has gained considerable attention because of the simplicity of building them and their of ease of parallelization (Berahas et al., 2019; Shi et al., 2023). In addition, a related research line is the simulation optimization via FD estimators (see, e.g., Fu, 2002), which is originated from Kiefer and Wolfowitz (1952). In this paper, we aim to establish an efficient FD estimator, servicing as a powerful tool for sensitivity analysis, applications in DFO, and so on.
FD methods encompass forward, backward, and central FD (CFD) methods. The primary challenge across all FD methods lies in determining the suitable perturbation. For example, increasing the perturbation simultaneously increases the bias and decreases the variance, making its selection crucial for minimizing the MSE ($=\mbox{bias}^{2}+\mbox{variance}$). Fox and Glynn (1989) point out that without employing common random numbers (CRNs), the optimal perturbation size of the CFD technique is $cn^{-1/6}$, corresponding to the optimal MSE of order $n^{-2/3}$, where $c$ is a constant, and $n$ is the number of sample pairs in the simulation. Although the order of the optimal perturbation corresponding to the various FD methods is known, the frontal constant, i.e., $c$, remains elusive.
The constant $c$ in particular is notoriously difficult to estimate, due to their reliance on additional model information, such as higher-order derivatives and simulation noises (Li and Lam, 2020). These quantities are typically unknown, and even estimating them is a more challenging task than the gradient estimation itself. Empirical evidence suggests that different constants may lead to substantially different MSEs.
In the vast literature on DFO and simulation optimization, where the FD estimator is used as a surrogate of the gradient, research on how to establish an asymptotically valid estimator of the optimal perturbation, and then the FD estimator itself, is rare. Typically, researchers tend to avoid estimating the constant $c$. Instead, in DFO literature, they find upper or lower bounds for higher-order derivatives and construct surrogates for these constants using these bounds (see, e.g., Shi et al., 2023). Meanwhile, in simulation optimization literature, only reasonable convergence rates of the perturbation and the FD estimator are guaranteed. Therefore, the former results in the FD estimator not converging to the true derivative, while the latter yields a suboptimal FD estimator (see, e.g., Fu, 2015).
In this paper, our aim is to establish an efficient FD estimator, in the sense of convergence to the true value with the optimal rate, sample-saving and practical implementation. Two crucial issues arise. The first is the estimation of $c$, which is fundamental for estimating the optimal perturbation. To estimate the constant consistently, the domain knowledge or prior knowledge, e.g., their upper or lower bounds, is insufficient. Therefore, some pilot samples need to be generated, and sample-driven methods should be proposed to estimate the constant based on these samples. However, since the ultimate goal is to obtain an FD estimator rather than a perturbation estimator, the number of pilot samples cannot be large. Thus, it has to be considered how to estimate the constants with relatively few samples. The second issue is establishing an efficient FD estimator. Given the constant estimator, the estimator of the optimal perturbation is obtained. A natural approach is to generate samples at the estimated optimal perturbation and average these samples to form the FD estimator. Intuitively, this estimator has the same asymptotic MSE as the traditional FD estimator. However, this approach is less efficient due to the under-utilization of samples, such as discarding pilot samples not generated at the estimated optimal perturbation. Therefore, effectively reusing these samples is crucial. In this paper, we propose novel methods to address these two issues.
1.1 Main Contributions
Specifically, the main contributions of this paper are as follows:
1. We propose a sample-driven method to estimate unknown constants and analyze their asymptotic properties. We estimate the constants primarily by leveraging the functional relationships between the bias and variance of the FD estimator and the perturbation. To achieve this, we propose a sample-driven method that combines the bootstrap technique with these functional relationships. Using pilot samples, we establish the bootstrap bias and variance of the FD estimator, proving that they maintain the functional relationships. Therefore, the constants are estimated based on the bootstrap bias and variance, which further lead to an estimator of the optimal perturbation. In addition, we establish the convergence rate of these estimators, serving as a guideline for generating the pilot samples.
2. We construct correlated samples at the estimated optimal perturbation to establish an efficient FD estimator and prove its theoretical properties. We innovatively transform pilot samples to resemble those generated at the estimated optimal perturbation, yielding correlated samples. Using these samples, we establish the efficient FD estimator, which we refer to as the correlation-induced FD estimator (Cor-FD). Although the correlation complicates theoretical analyses, it surprisingly results in the proposed estimator having a control variate configuration, thereby reducing variance and, in some scenarios, decreasing bias compared to the traditional optimal FD estimator. Notably, the total samples can be used as pilot samples, which leads to the estimator’s potential availability even in cases of small sample sizes.
3. We establish algorithms for the proposed estimator’s practical implementation and application in solving DFO problems, demonstrating its practical benefits through numerical experiments. Theoretical properties provide a guiding framework for practical implementation. Based on this, we propose a detailed algorithm to derive the proposed estimator. Furthermore, we apply the estimator in various numerical examples. Numerical results are consistent with the theoretical analysis, and illustrate that both the perturbation estimator and the FD estimator preform well compared to other approaches, particularly with small sample sizes. Additionally, by integrating the proposed Cor-FD estimator into a quasi-Newton algorithm, we develop an algorithm to solve DFO problems. This algorithm is compared numerically with some popular algorithms, highlighting the advantages of our algorithm over others.
1.2 Literature Review
In stochastic gradient estimation, there are direct and indirect approaches (Fu, 2015). The former includes infinitesimal perturbation analysis (Ho et al., 1983; Heidelberger et al., 1988), likelihood ratio or score function (Glynn, 1990; Rubinstein, 1986; Reiman and Weiss, 1989), weak differentiation (Heidergott et al., 2006, 2010) and generalized likelihood ratio method (Peng et al., 2018). These gradient estimation methods need to understand the underlying stochastics of the system. However, as systems become increasingly complex, it is often easier to experiment with them than to fully understand them (Golovin et al., 2017). In such cases, indirect approaches, including the FD estimator we study in this paper, are more suitable.
The indirect gradient estimation means that the gradient is estimated solely from noisy function evaluations. These approaches include traditional FD (see, e.g., Chapter 7 in Glasserman, 2013)(Chapter 7 in Asmussen and Glynn, 2007; L’Ecuyer, 1991; Fu, 2015), Gaussian smoothing FD (Nesterov and Spokoiny, 2015), random FD (RFD) (see, e.g., Salimans et al., 2017; Shamir, 2017; Fazel et al., 2018), etc. In these methods, all estimators are established via the average of independent samples, and the crucial aspect is determining the perturbation size (also known as the differencing interval in some literature like Berahas et al. (2019)). Typically, as the perturbation decreases, the variance of the estimator increases while the bias decreases. To select a suitable perturbation for the traditional FD estimator, Fox and Glynn (1989) and Zazanis and Suri (1993) tradeoff its bias and variance by minimizing its MSE, and then provide the optimal orders of the perturbation. However, constants in the perturbation are typically unknown, making the FD estimator optimal only in the sense of convergence rate. In contrast, our work proposes an FD estimator based on correlated samples, which not only achieves the same MSE convergence rate as the optimal traditional FD estimator but also exhibits variance reduction compared to the latter.
Recently, to circumvent the estimation of the optimal perturbation, Lam et al. (2023) propose minimizing the worst-case asymptotic ratio of the MSE of a proposed estimator to that of the traditional FD estimator w.r.t. model characteristics, including the perturbation. In this framework, a general weighted FD estimator is constructed by introducing a hyperparameter. In contrast, our work attempts to establish an FD estimator by directly estimating the optimal perturbation first.
The closest work to ours is Li and Lam (2020), which suggest a two-stage technique. In the first stage, pilot samples are used to estimate the optimal perturbation; in the second stage, additional samples are generated according to this estimated perturbation to construct an FD estimator. Although this two-stage technique yields an FD estimator with asymptotically optimal MSE, the under-utilization of samples, such as discarding the pilot samples, results in reduced efficiency. In our work, we propose a sample-driven method using the bootstrap technique based on some pilot samples inspired by by Zhang et al. (2022) to estimate the optimal perturbation. Then we indicate that all samples for establishing our proposed estimator can be used as pilot samples by transforming them to those at the estimated optimal perturbation. This twice sample-reusing procedure not only enhances the efficiency of samples but also potentially allows our proposed estimator to be practically implemented even with a small sample size.
An important study of FD approximation is its application in stochastic optimization (Chau and Fu, 2014; Fu, 2015). Kiefer and Wolfowitz (1952) introduce the earliest FD-based stochastic approximation (KW) algorithm, where the step lengths and perturbations must be carefully chosen. To improve the KW algorithm, Broadie et al. (2011) provide a scaled-and-shifted (for the step length and perturbation, respectively) KW algorithm. Recently, Chen et al. (2023) apply this algorithm to solve an optimization problem in a queueing system. In multivariate settings, Spall (1992, 1997) propose the simultaneous perturbation stochastic approximation (SPSA) algorithm. Recently, Xu and Zheng (2023) propose a KW-type algorithm using the RFD estimator for solving multidimensional simulation optimization problems. Hu and Fu (2024) demonstrate that to ensure the convergence of these algorithms, the rate of convergence of step lengths to zero must be faster than the increase in the variance of the FD estimator caused by the decrease in perturbations.
Another research line of stochastic optimization via FD approximation is the so called DFO area (Berahas et al., 2019; Larson et al., 2019; Shi et al., 2023), where the step lengths are not required to tend to 0, and the bounded error of the FD estimator is typically assumed, so the algorithms’ results only converge to a neighborhood of the optimums. In contrast, the algorithm we establish for solve DFO problems is based on our proposed CFD estimator. To avoid digression, we only provide numerical results, rather than theoretical analyses, to demonstrate the algorithm’s good performance, without assuming the convergence of step lengths to zero.
1.3 Organization
@@
The rest of this paper is organized as follows. Section 2 gives some background on using the finite difference approximation for gradient estimation. In Section 3, we propose a sample-driven method, apply it to perturbation estimation, and provide theoretical guarantees for it. Section 4 constructs correlated samples for establishing the efficient FD estimator, and provides theoretical evidence supporting its effectiveness. For practical implementation, a detailed algorithm is proposed in Section 5, and we verify the theoretical results through numerical experiments in Section 6. Section 7 applies our proposed FD estimator to solve DFO problems, followed by conclusions in Section 8. Proofs, discussions and additional results are provided in the appendix.
@@ 2 Background
In this section, we introduce the background of the finite difference method for stochastic gradient estimation. More details can be found in Asmussen and Glynn (2007), Fu (2006) and Glasserman (2013). Consider the models that depend on a single parameter $\theta$, where $\theta$ varies within some range [^4] When $\Theta\in\mathbb{R}^{d},d>1$, we can repeat the same operation for each direction. Therefore, without loss of generality, we assume $\Theta\in\mathbb{R}$. $\Theta\subset\mathbb{R}$. Denote that $\alpha(\cdot)$ is a performance measure of interest and assume that it can only be estimated by simulation. Within the simulation for any chosen $\theta\in\Theta$, each trail gives an unbiased but noisy estimator of $\alpha(\theta)$, denoted by $Y(\theta)$, i.e., $\alpha(\theta)=\mathbb{E}[Y(\theta)]$. Suppose that we do not use the CRNs, i.e., for $\theta_{1}\neq\theta_{2}\in\Theta$, $Y(\theta_{1})$ and $Y(\theta_{2})$ are independent. We would like to estimate the first-order derivative $\alpha^{\prime}(\theta_{0})$, where $\theta_{0}\in\Theta$ is the point of interest. To estimate $\alpha^{\prime}(\theta_{0})$, we prefer the CFD method over the forward (backward) FD method because compared to the latter, the CFD method yields an improvement in the convergence rate of the bias. It is worth mentioning that the methods in Sections 3 and 4 are seamlessly adapted to the forward (backward) FD estimator.
@@
Specifically, the CFD scheme sets the perturbation parameter $h>0$. At $\theta_{0}+h$, the simulation is repeated independently, and $n$ independently and identically distributed (i.i.d.) observations $\{Y_{1}(\theta_{0}+h),...,Y_{n}(\theta_{0}+h)\}$ are obtained. Also, $n$ i.i.d. observations $\{Y_{1}(\theta_{0}-h),...,Y_{n}(\theta_{0}-h)\}$ are simulated at $\theta_{0}-h$. Using the sample average, we construct a CFD estimator of $\alpha^{\prime}(\theta_{0})$, denoted by
$\widehat{\Delta}_{n,h}=\frac{1}{n}\sum_{i=1}^{n}\Delta_{i}(h)=\frac{1}{n}\sum_{i=1}^{n}\frac{Y_{i}(\theta_{0}+h)-Y_{i}(\theta_{0}-h)}{2h}.$ (3)
For any $i=1,...,n$, $Y_{i}(\theta_{0}+h)$ and $Y_{i}(\theta_{0}-h)$ are the unbiased estimators of $\alpha(\theta_{0}+h)$ and $\alpha(\theta_{0}-h)$, respectively, so $\Delta_{i}(h)$ and $\widehat{\Delta}_{n,h}$ are both the unbiased estimators of $\widetilde{\alpha}_{h}^{\prime}(\theta_{0})$. Let $Z_{i}(h)$ denote the zero-mean error associated with $\Delta_{i}(h)$, and let $\bar{Z}_{n,h}=\sum_{i=1}^{n}Z_{i}(h)/n$ be the zero-mean error associated with $\widehat{\Delta}_{n,h}$. Then,
@@
The choice of $h$ is crucial when constructing $\widehat{\Delta}_{n,h}$. If $h$ is too large, the bias of $\widehat{\Delta}_{n,h}$ will be large, else if $h$ is too small, the variance of $\widehat{\Delta}_{n,h}$ will explode. A proper $h$ should balance the squared bias and variance of $\widehat{\Delta}_{n,h}$. To calculate the bias and variance elaborately, we make the following assumptions.
@@
The standard deviation of $Y(\theta)$, denoted by $\sigma(\theta)$, is continuous at $\theta_{0}$ and positive.
@@
Subtraction eliminates $\alpha^{(2)}(\theta_{0})$, leaving $\widetilde{\alpha}_{h}^{\prime}(\theta_{0})=\alpha^{\prime}(\theta_{0})+Bh^{2}+o(h^{2})$, where $B=\alpha^{(3)}(\theta_{0})/6$. Therefore, the bias of $\widehat{\Delta}_{n,h}$ is
@@
It follows from (7) that $h$ controls a trade-off between the bias and variance. Minimizing the MSE w.r.t. $h$ leads to the optimal perturbation
@@
Thus the optimal bias, variance and MSE are
@@
Theoretically, $h^{*}$ in (8) yields an optimal $\widehat{\Delta}_{n,h^{*}}$. Although we know that $h^{*}=Cn^{-1/6}$ from (8), where $C=\left(\sigma^{2}(\theta_{0})/(4B^{2})\right)^{1/6}$, the constant $C$ is typically unknown. The definition of $C$ indicates that its estimation depends on the model information, such as $\alpha^{(3)}(\theta_{0})$ and $\sigma^{2}(\theta_{0})$. However, estimating these values is more challenging than estimating the gradient itself, because it involves the higher-order derivatives of $\alpha(\theta)$. Furthermore, experimental findings indicate that assigning a perturbation with a significant deviation from the true one leads to substantial discrepancies in the MSE of the CFD estimator. These motivate us to explore an approach to estimate $h^{*}$.
3 A Sample-driven Method for Constant Estimation
In this section, we propose a sample-driven method to estimate $C$, specifically $B$ and $\sigma^{2}(\theta_{0})$. Our inspiration comes from (5) and (6), which demonstrate that $\mathbb{E}\widehat{\Delta}_{n,h}$ and $\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]$ are linear w.r.t. $h^{2}$ and $1/h^{2}$, respectively. If we can accurately calculate $\mathbb{E}\widehat{\Delta}_{n,h}$ and $\mathrm{Var}\left[\widehat{\Delta}_{n,h}\right]$ for any given $h$, then choosing $h_{1}\neq h_{2}$ and solving the system of linear equations